# A coupled-trajectory approach for decoherence, frustrated hops and internal consistency in surface hopping

### v2

@@ Abstract
    $\Delta E^{\alpha}_{\text{Qmom}}=\Delta E^{\alpha}_{\text{hop}}\frac{\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}+\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}$ (4)
@@
In general, in CCT-TSH, we assume that the kinetic energy of the hopping trajectory $\alpha$ is used first, before requesting the cooperation of the surrounding trajectories via the energy sharing. Therefore, the velocity rescaling options used in the equity-based and in the overlap-based sharing schemes are either isotropic, meaning that all components of the velocity vector are equally rescaled, or hybrid, meaning that the component of the velocity along the nonadiabatic coupling vector is rescaled if possible, and otherwise it is isotropic. Nonetheless, we will show that the velocity rescaling procedure does not impact the results.
@@
It is worth discussing here the decoherence problem of surface hopping and its relation to the internal consistency. Internal consistency refers to the agreement between the two ways of estimating the population of the electronic states in surface hopping: the fraction of trajectories associated to, or evolving in, an electronic state (F); the trajectory-averaged population propagated along each trajectory (P), since usually a quantum-dynamical evolution equation for the electronic state is integrated along each trajectory. However, in general and unless decoherence corrections are used, it cannot be ensured that the two estimates agree with each other. The original hopping algorithm [2] ensures this internal consistency by using an average hopping probability over the swarm of trajectories for which the two estimates for the populations are the same. However, as soon as an independent trajectory formalism is invoked and the hopping probability is calculated at the level of a single trajectory, internal consistency is no longer guaranteed. The coupled-trajectory schemes that we propose here are naturally designed to reintroduce the idea of an average hopping probability to restore internal consistency. However, within CT-TSH this might introduce unphysical artifacts. Such an average hopping algorithm will assign the same hopping probability to all trajectories independently of their position. While this might ensure a correct distribution of the wavepacket/swarm with respect to the electronic states, it neglects the locality of the nonadiabatic couplings and the spatial distribution of the wavepacket, causing trajectories to undergo hops even in regions of negligible nonadiabatic coupling. This limitation has previously been overcome through the consensus surface hopping algorithm.[35] For CCT-TSH, we propose, by contrast, to forgo the stochastic hopping algorithm and introduce a deterministic hopping. [49] Since the electronic evolution is derived from the EF equation for the conditional amplitude, we drive the trajectory hops based on the values of the electronic populations predicted by this equation. In a way, the nuclear trajectories in CCT-TSH can be regarded merely as a support on which the time-dependent electronic state and the quantum momentum are evaluated. Therefore, it is crucial that the trajectories evolve following the electronic populations. For this reason, in CCT-TSH, the trajectories evolve in the electronic state with the largest population.
@@
The CCT-TSH algorithms illustrated above are tested on the 30-dimensional linear vibronic coupling (LVC) model of fulvene.[29] Upon photoexcitation to the first excited state S1, fulvene decays through two conical intersections, the first one is a sloped conical intersection seam and is associated to a stretching of the ethenylic bond and the second one is a peaked conical intersection and is associated to the twist of the methylene group [32, 50]. This LVC model of fulvene reproduces only the sloped conical intersection, but not the peaked intersection. However, the latter appears to be not important in the simulated short time dynamics as argued previously [51, 29]. Furthermore, we are interested in correctly capturing the challenging dynamics at the slope conical intersection, where the wavepacket/swarm is partially reflected and thus crosses multiple times the nonadiabatic coupling region. Also, note that the molecular system of fulvene has been shown to be highly sensitive to the parameters of the surface hopping algorithm, in terms of procedure to treat frustrated hops, velocity rescaling and time step.[32, 33, 31] Therefore, it is a suitable test system to verify the robustness of the proposed CCT-TSH schemes. All simulations have been initialized with a swarm of 500 trajectories with positions and momenta Wigner sampled from the vibrational ground state of the electronic LVC ground state potential. The trajectories were propagated with a time step of 0.1 a.t.u.
@@
Figure 2: Adiabatic population decays of S1 predicted by the different surface hopping schemes discussed in the main text. The lighter lines in each panel show the dynamics with different velocity rescaling procedures and treatments of frustrated hops, while the bold darker lines show the average of the individual runs. Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
@@
In Fig. 2, we show the adiabatic population curves of the fulvene LVC model upon excitation to the S1 state. For the standard surface hopping approaches, TSH, TSH-ED, CT-TSH, six different dynamics are averaged, using three different schemes to rescale the velocity after a hop, with each option used in combination with velocity inversion or not in the case of frustrated hops. The three schemes for the velocity rescaling are (1) isotropic, (2) along the nonadiabatic coupling vectors, or (3) mixed, meaning that if (2) is not possible then (1). In a standard application of any surface hopping scheme to a molecular problem, each of these approaches would per se be regarded as a justified choice, which is why we chose to use the repetitions with different schemes to give us some information on the uncertainty of the results. It can be clearly seen that for both TSH and CT-TSH internal consistency is not ensured, and in both cases, the electronic populations (P) decay slower than the fractions of trajectories (F). TSH-ED restores, as expected, almost perfect internal consistency. The population dynamics are very similarly predicted by all three methods, with a sharp decrease of S1 population occurring after 250 a.t.u., but since the wavepacket/swarm of trajectories is partially reflected, we observe stepwise decays to around only 10% of S1 population at the end of the simulated dynamics. When looking at CCT-TSH, it can be observed that all three proposed schemes restore closely internal consistency. The population decay predicted using the equity-based and overlap-based energy sharing schemes are almost identical, which points to an encouraging robustness of the algorithm. However, the quantum-momentum-based sharing scheme yields a different population decay with more pronounced coherent oscillations of the population. Given that the quantum-momentum-based energy sharing acts at every surface hop and not only when a hop would be frustrated on the single trajectory level, this result is hardly surprising.
@@
The number of frustrated hops during the dynamics varies strongly between the different methods (see Table 1). As expected, when we allow only the component of the velocity along the nonadiabatic coupling vectors to be rescaled, the number of frustrated hops is quite large in all methods. Applying a decoherence correction as in TSH-ED, not only improves the internal consistency but also strongly decreases the number of frustrated hops. If the full velocity vector can be rescaled isotropically, there are no (or a negligible number of) frustrated hops during the dynamics. The number of frustrated hops during the CT-TSH dynamics is comparable to TSH, but surprisingly much larger if only the component of the velocity along the nonadiabatic coupling vectors can be rescaled. This issue is completely alleviated, albeit by construction, when using the equity-based and overlap-based CCT-TSH schemes. By contrast, the energy sharing scheme based on the quantum momentum shows a large number of frustrated hops. This can be explained by the modified hopping scheme where at each time step the energy sharing is attempted.
@@
In order to ascertain that the energy sharing scheme does not lead to artifacts in the spatial distribution of the wavepacket, we report the distribution of the trajectories along the normal mode corresponding to the C=CH2 stretching. This mode was chosen based on previous works where it has been observed that the main sloped conical intersection between S1 and S0 is accessed via a stretching of the C=CH2 bond. [32, 52] Each trajectory was convoluted with a Gaussian of width 0.01 along this mode. We note that in Fig. 3 the $y$-axis is given in mass-weighted normal mode coordinates. Negative values correspond to an elongation of the bond, while positive values correspond to a contraction. The dynamics of TSH, TSH-ED and CT-TSH were averaged over all the six different runs. Over time, the swarm of trajectories shows harmonic oscillations along this bond, initially elongating it, which drives the population decay. After the first elongation, around 500 to 1000 a.t.u., a splitting of the swarm of trajectories can be observed. One part contracts the bond further whereas another part contracts only until the equilibrium position before elongating again. This splitting can be attributed to the electronic population, where the part that decays to the ground state harmonically oscillates between $-4$ and $4$ and the part that gets reflected on the excited state oscillates only in the elongation between 00 and $-4$. In all methods, TSH, TSH-ED and CT-TSH, these two branches are clearly distinguishable between 500 and 2000 a.t.u. A very similar behavior is exhibited by the CCT-TSH trajectories within the equity-based and overlap-based schemes. As with the population traces, their behavior seems to be in excellent agreement with TSH-ED results. The rescaling proportional to the quantum momentum shows a similar overall behavior, however the branching is less clear as the trajectories are overall more spatially delocalized.
@@
Figure 3: Spatial evolution of the swarms of trajectories along the C=CH2 stretching mode. The $y$-axis is given in mass-weighted normal mode coordinates, negative values correspond to a stretching of the C=CH2 bond, positive values to a contraction. TSH, TSH-ED and CT-TSH are averaged over the six different dynamics.
@@
The kinetic energy along this normal mode (shown in Fig. 4) shows that the redistribution of kinetic energy in equity-based and overlap-based CCT-TSH does not alter significantly the average kinetic energy along that normal mode. For TSH, TSH-ED and CT-TSH, the six different runs are shown. The average kinetic energies of all methods are in good agreement and CCT-TSH yields values that lie within the uncertainty spread of the different standard surface hopping schemes. Through the kinetic energy, it becomes clear that the quantum-momentum-based approach does not yield reliable results, as it largely increases the kinetic energy in this mode.
@@
Figure 4: Average kinetic energy of the ensemble of trajectories along the C=CH2 stretching mode.

### v3

@@ 2 Coupled-trajectory methods combined with surface hopping
    $\Delta E^{\alpha}_{\text{Qmom}}=\Delta E^{\alpha}_{\text{hop}}\frac{|\boldsymbol{q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}|}{|\boldsymbol{v}^{\alpha}\cdot\boldsymbol{d}_{IJ}^{\alpha}|+|\boldsymbol{q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}|}$ (15)
@@
In general, in CCT-TSH, we assume that the kinetic energy of the hopping trajectory $\alpha$ is fully used first, before requesting the cooperation of the surrounding trajectories via the energy sharing. Therefore, the velocity rescaling options used in the equity-based and in the overlap-based sharing schemes are either isotropic, meaning that all components of the velocity vector are equally rescaled, or hybrid, meaning that the component of the velocity along the NACV is rescaled if possible, and otherwise it is isotropic. Nonetheless, we will show that the velocity rescaling procedure does not impact the results. As indicated above, in all energy-sharing options, the contributing trajectories are selected based on their kinetic energy, which needs to be above a certain threshold. This option is considered mainly to avoid that such contributing trajectories suffer later on in the dynamics from the problem of frustrated hops.
@@
It is worth discussing at this point the decoherence problem of surface hopping 24, 16 and its relation to the internal consistency. Internal consistency refers to the agreement between the two ways of estimating the population of the electronic states in surface hopping: the fraction of trajectories associated to, or evolving in, an electronic state (F); the trajectory-averaged population propagated along each trajectory (P), since usually a quantum-dynamical evolution equation for the electronic state is integrated along each trajectory. However, in general and unless decoherence corrections are used, it cannot be ensured that the two estimates agree with each other. The original hopping algorithm 2 ensures this internal consistency by using an average hopping probability over the swarm of trajectories for which the two estimates for the populations are the same. However, as soon as an independent trajectory formalism is invoked and the hopping probability is calculated at the level of a single trajectory, internal consistency is no longer guaranteed. The coupled-trajectory schemes that we propose here are naturally designed to reintroduce the idea of an average hopping probability to restore internal consistency. However, within CT-TSH this might introduce unphysical artifacts. Such an average hopping algorithm will assign the same hopping probability to all trajectories regardless of their position. While this might ensure a correct distribution of the wavepacket/swarm with respect to the electronic states, it neglects the locality of the nonadiabatic couplings and the spatial distribution of the wavepacket, causing trajectories to undergo hops even in regions of negligible nonadiabatic coupling. This limitation has previously been overcome through the CSH algorithm, 36 where an ensemble of stochastic trajectories is used to solve the semiclassical Liouville equation. 77, 78, 79, 80, 81, 82, 83, 84, 85, 86 Note that in CSH, the velocity rescaling procedure is circumvented, since the total energy is naturally conserved over the entire ensemble of trajectories thanks to the underlying structure based on the Liouville equation. Therefore, while CCT-TSH might seem similar to CSH, it is fundamentally a different approach.
For CCT-TSH, we propose an alternative approach for the hopping algorithm, namely to forgo the stochastic scheme and to introduce a deterministic hopping, as also done in MASH. 50 Since the electronic evolution is derived from the EF equation for the conditional amplitude (4), we drive the trajectory hops based on the values of the electronic populations predicted by this equation. In a way, the nuclear trajectories in CCT-TSH can be regarded merely as a support on which the time-dependent electronic state and the quantum momentum are evaluated. Therefore, it is crucial that the trajectories evolve following the electronic populations. For this reason, in CCT-TSH, the trajectories evolve in the electronic state with the largest population. Nonetheless, we will assess the performance of this “deterministic” version of CCT-TSH by comparing it to its fewest-switches version, where the active state is, instead, chosen as in CT-TSH.
3 Fulvene and DMABN dynamics with coupled trajectories
The CCT-TSH algorithms illustrated above are tested on the 30-dimensional linear vibronic coupling (LVC) model of fulvene and on the 57-dimensional LVC model of DMABN.29
@@ 3 Fulvene and DMABN dynamics with coupled trajectories
Upon photoexcitation to the first excited state S1, fulvene decays through two conical intersections, the first one is a sloped conical intersection seam and is associated to a stretching of the ethenylic bond and the second one is a peaked conical intersection and is associated to the twist of the methylene group 32, 87. This LVC model of fulvene reproduces only the sloped conical intersection, but not the peaked intersection. However, the latter does not appear to be important in the simulated short-time dynamics as argued previously 88, 29. Furthermore, we are interested in correctly capturing the challenging dynamics at the slope conical intersection, where the wavepacket/swarm is partially reflected and thus crosses multiple times the nonadiabatic coupling region. Also, note that the molecular system of fulvene has been shown to be highly sensitive to the parameters of the surface hopping algorithm, in terms of procedure to treat frustrated hops, velocity rescaling and time step.32, 33, 31, 35 Therefore, it is a suitable test system to verify the robustness of the proposed CCT-TSH schemes. Concerning DMABN, the electronic structure and photodynamics of DMABN have been previously studied quite extensively in Refs. [89, 90, 91, 92, 32, 93, 94]. For the LVC model proposed in Ref .[93], it was found that DMABN encounters an “immediate” conical intersection between states S1 and S2 after photoexcitation to S2, which is accessible directly by the photoexcited wavepacket located in the Franck-Condon region. This leads to a fast population transfer from state S2 to S1. Ab initio studies have shown that DMABN has an additional conical intersection between S2 and S1 that is associated with a rotation of the dimethylamino group 89, but the LVC model does not account for this conical intersection.
3.1 Computational details
All algorithms employed for the simulations performed here are implemented in G-CTMQC 95, which is available open-source in GitLab.
All simulations have been initialized with a swarm of 500 trajectories with positions and momenta Wigner sampled from the vibrational ground state of the electronic LVC ground state potential centered at the origin (in both position and momentum). The trajectories were propagated with a time step of 0.1 a.t.u. for 4200 a.t.u (about 100 fs of dynamics are simulated in this way). Such a small integration time step is necessary because the multiple time step procedure is currently being tested in G-CTMQC and it is, thus, not used in the calculations performed here. Some simulations have also been run by Wigner-sampling only the positions while initializing all the trajectories with the same zero momentum to assess the robustness of the algorithms with different sets of initial conditions. [^2] Note that we expect the final results to be affected upon changing the initial conditions, since the solution of the ordinary differential equations representing the algorithms clearly depends on the initial conditions. However, our tests aim to assess the robustness of the energy-sharing procedures within different ways of initializing the dynamics.
Numerical tests are performed using the original surface hopping algorithm (TSH) and the modified version that includes decoherence using the energy-based approach (TSH-ED) proposed by Granucci and Persico 96. TSH and TSH-ED are based on Eqs. (8) and (9) combined with the fewest-switches procedure to select randomly the active state. In TSH-ED, the standard value of $0.1$ Ha is used for the energy parameter in the exponential decay applied to the electronic coefficients that are not associated to the active state.
G-CTMQC allows to perform surface-hopping-based calculations using different schemes:
  • for the treatment of frustrated hops (w/o inversion of the nuclear velocity for the trajectory experiencing a frustrated hop),
  • for the rescaling of the velocity after a hop (isotropical, along the NACVs, or a mixture of the two to avoid frustrated hops),
  • for the hopping procedure (either using fewest switches or by forcing the hop towards a state whose population reaches a certain threshold, which is chosen to be 0.5 in our calculations),
  • for the energy sharing (equity-based, overlap-based, quantum-momentum-based).
We provide as Supporting Information a script to help generate the input files to run all the calculations reported below, thus, all our data are easily reproducible.
G-CTQMC is interfaced with QuantumModelLib 97, a library of analytical potentials that performs the diagonalization of the electronic LVC Hamiltonian
$\hat{H}_{el}(\mathbf{q})=\hat{H}^{\mathrm{LVC}}(\mathbf{q})-\left(\sum_{n=1}^{N_{\mathrm{modes}}}-\frac{\omega_{n}}{2}\frac{\partial^{2}}{\partial q_{n}^{2}}\right)\hat{I}_{N_{\mathrm{st}}\times N_{\mathrm{st}}},$ (16)
to provide analytically adiabatic energies and gradients as well as the NACVs. Here, the second term on the right-hand side is the nuclear kinetic energy operator and the first term is the full LVC Hamiltonian 98, 99, 100, 101, 102, 103, 104
$\hat{H}^{\mathrm{LVC}}(\mathbf{q})=\hat{H}^{(0)}(\mathbf{q})+\hat{W}^{0}(\mathbf{q})+\hat{W}^{1}(\mathbf{q}),$ (17)
that has an analytical form, which (i) approximates harmonically the coupled (diabatic) potential energy surfaces in the vicinity of the Franck-Condon region (ii) including linear terms and (iii) coupling linearly the potentials associated to different electronic states. The LVC Hamiltonian is constructed in the diabatic representation, thus the electronic Hamiltonian is not diagonal in this basis, and the nuclear modes used are the normal modes $\mathbf{q}$ identified from the harmonic analysis of the ground-state potential energy surface in the Franck-Condon region. The zero-th order Hamiltonian is diagonal in the space of the electronic states
$\hat{H}^{(0)}(\mathbf{q})=\left(\sum_{n=1}^{N_{modes}}-\frac{\omega_{n}}{2}\frac{\partial^{2}}{\partial q_{n}^{2}}+\frac{\omega_{n}}{2}q_{n}^{2}\right)\hat{I}_{N_{\mathrm{st}}\times N_{\mathrm{st}}},$ (18)
with mass-frequency weighted normal coordinates $\mathbf{q}=\{q_{n}\}_{n=1,\ldots,N_{modes}}$ and harmonic frequencies $\omega_{n}$. For $N_{\mathrm{st}}$ electronic states, $\hat{W}^{0}(\mathbf{q})$ and $\hat{W}^{1}(\mathbf{q})$ are $N_{\mathrm{st}}\times N_{\mathrm{st}}$ matrices with elements
${W}^{0}_{lm}=E_{l}\delta_{lm}$ (19)
and
${W}^{1}_{lm}(\mathbf{q})=\begin{cases}\sum_{n}k^{(l)}_{n}q_{n}&l=m\\ \sum_{n}\lambda^{(l,m)}_{n}q_{n}&l\neq m\end{cases}$ (20)
The parameters $E_{l}$, $k^{(l)}_{n}$, $\lambda^{(l,m)}_{n}$, and $\omega_{n}$ used in our simulations are reported in Ref. [93].
3.2 Numerical results
@@
Figure 2: Adiabatic population decays of S1 in fulvene predicted by the different surface hopping schemes discussed in the main text. The lighter lines in each panel show the dynamics with different velocity rescaling procedures and treatments of frustrated hops, while the bold darker lines show the average of the individual runs. Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
@@
In Fig. 2, we show the adiabatic population curves of the fulvene LVC model upon excitation to the S1 state. For the standard surface hopping approaches, namely for TSH, TSH-ED, CT-TSH, six different dynamics are averaged, using three different schemes to rescale the velocity after a hop, with each option used in combination with velocity inversion or not in the case of frustrated hops. The three schemes for the velocity rescaling are (1) isotropic, (2) along the NACVs, or (3) mixed, meaning that if (2) is not possible then (1). In a standard application of any surface hopping scheme to a molecular problem, each of these approaches would per se be regarded as a justified choice, which is why we chose to use the repetitions with different schemes to give us some information on the uncertainty of the results. In the case of CCT-TSH, in Fig. 2, we compare the results of the three energy-sharing schemes, i.e., equity-based, overlap-based, quantum-momentum-based, employed with the option of rescaling the velocity along the NACV and inverting the velocity in the case of frustrated hops. Note that using other velocity rescaling options does not change dramatically the results. In addition, we use a deterministic selection of the active state, depending on the state with largest populations.
Note that in Fig. 2 we aim to point out the quite large distributions of values of the S1 population, especially in TSH and CT-TSH, rather than the actual calculated values, which is why we only distinguish clearly the continuous lines – electronic populations (P) determined from the electronic evolution equation – and the dashed lines – fractions of trajectories (F) evolving in S1. The thick continuous and dashed lines, green for TSH, blue for TSH-ED and purple for CT-TSH, are the averages of the corresponding populations, with the spreading of the light-colored lines providing an idea of the associated uncertainty.
It can be clearly seen that for both TSH and CT-TSH internal consistency is not ensured, and in both cases, the electronic populations (P) decay slower than the fractions of trajectories (F). TSH-ED restores, as expected, almost perfect internal consistency. The population dynamics are very similarly predicted by all three methods, with a sharp decrease of S1 population occurring after 250 a.t.u., but since the wavepacket/swarm of trajectories is partially reflected, we observe stepwise decays to around only 10% of S1 population at the end of the simulated dynamics. When looking at CCT-TSH, it can be observed that all three proposed schemes restore closely internal consistency and the population decays are almost identical with the three energy sharing schemes, which points to an encouraging robustness of the algorithm. However, the quantum-momentum-based sharing scheme yields a large number of frustrated hops (see Table 1). This result is hardly surprising for two reasons: the approximate nature of the quantum momentum, discussed above; the quantum-momentum-based energy sharing acts at every surface hop and not only when a hop would be frustrated on the single trajectory level. Note, in addition, that while the quantum-momentum-based energy sharing might seem the most physically appropriate, the choice to partition the energy according to Eqs. (14) and (15) can be considered quite arbitrary and certainly different partitioning schemes are possible. For instance, in a case where the potential energy surfaces associated to two or more (coupled) electronic states remain parallel to each other (see Refs. [105, 106] for discussions on the importance of these situations), Eq. (15) would yield always zero, since the difference in gradients accumulated over time is zero as well. Similarly, the equity-based scheme might become “unphysical” if two or more branches of the nuclear density/trajectories travel far away from each other, thus, one would not expect plausible that all branches contribute to hops occurring locally in a region of a single branch.
@@
The number of frustrated hops during the dynamics varies strongly between the different methods (see Table 1). As expected, when we allow only the component of the velocity along the NACVs to be rescaled, the number of frustrated hops is quite large in all methods. Applying a decoherence correction as in TSH-ED, not only improves the internal consistency but also strongly decreases the number of frustrated hops. If the full velocity vector can be rescaled isotropically, there are no (or a negligible number of) frustrated hops during the dynamics. The number of frustrated hops during the CT-TSH dynamics is comparable to TSH, but surprisingly much larger if only the component of the velocity along the NACVs can be rescaled. This issue is completely alleviated, albeit by construction, when using the equity-based and overlap-based CCT-TSH schemes. By contrast, the energy sharing scheme based on the quantum momentum shows a large number of frustrated hops. This can be explained by the modified hopping scheme where at each time step the energy sharing is attempted.
@@
Figure 3: Spatial evolution of the swarms of trajectories along the C=CH2 stretching mode in fulvene. The $y$-axis is given in mass-weighted normal mode coordinates, negative values correspond to a stretching of the C=CH2 bond, positive values to a contraction. TSH, TSH-ED and CT-TSH are averaged over the six different dynamics.
In order to ascertain that the energy sharing scheme does not lead to artifacts in the spatial distribution of the wavepacket, we report the distribution of the trajectories along the normal mode corresponding to the C=CH2 stretching. This mode was chosen based on previous works where it has been observed that the sloped conical intersection between S1 and S0 is accessed via a stretching of the C=CH2 bond. 87, 93 Each trajectory was convoluted with a Gaussian of width 0.01 along this mode. We note that in Fig. 3 the $y$-axis is given in mass-weighted normal mode coordinates. Negative values correspond to an elongation of the bond, while positive values correspond to a contraction. The dynamics of TSH, TSH-ED and CT-TSH were averaged over all the six different runs. Over time, the swarm of trajectories shows harmonic oscillations along this bond, initially elongating it, which drives the population decay. After the first elongation, around 500 to 1000 a.t.u., a splitting of the swarm of trajectories can be observed. One part contracts the bond further whereas another part contracts only until the equilibrium position before elongating again. This splitting can be attributed to the electronic population, where the part that decays to the ground state harmonically oscillates between 4-4 and $4$ and the part that gets reflected on the excited state oscillates only in the elongation between 0 and 4-4. In all methods, TSH, TSH-ED and CT-TSH, these two branches are clearly distinguishable between 500 and 2000 a.t.u. A similar behavior is exhibited by the CCT-TSH trajectories, and as with the population traces, their behavior seems to be in excellent agreement with TSH-ED results.
@@
The kinetic energy along this normal mode (shown in Fig. 4) shows that the redistribution of kinetic energy in all flavors of CCT-TSH does not alter significantly the average kinetic energy along that normal mode. For TSH, TSH-ED and CT-TSH, the six different runs are shown, while for CCT-TSH we show, as in Fig. 2, the results obtained by reflecting the velocity of the trajectory that undergoes a frustrated hop (even though in the equity-based and overlap-based versions there are no frustrated hops) and by rescaling the velocity along the NACV. The average kinetic energies of all methods are in good agreement and CCT-TSH yields values that lie within the uncertainty spread of the different standard TSH-ED schemes.
@@
Figure 4: Average kinetic energy of the ensemble of trajectories along the C=CH2 stretching mode in fulvene.
Having assessed the good performance and the stability of the energy-sharing schemes in fulvene on the excited-state population and on some aspects of the nuclear dynamics, we report additional comparisons of numerical schemes for CCT-TSH aiming to demonstrate its robustness.
Figure 5: Top panel: Population of S1 in fulvene as a function of time calculated with CCT-TSH using different flavors of the overlap-based: using the fewest-switches procedure combined with a Wigner sampling of the initial conditions (in light-brown), using the fewest-switches procedure but sampling harmonically only the positions while setting all momenta of the trajectories at time $t=0$ to be equal to $\mathbf{P}_{\nu}=\mathbf{P}_{0}=\mathbf{0}$ $\forall\nu$ (in yellow), using the deterministic selection of the active state combined with a Wigner sampling (in orange, as in Fig. 2). Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F). Bottom panel: Population of the diabatic state initially populated compared to MCTDH (in black) calculations of Ref. [29]
In Fig. 5, we show the populations of S1 (top panel) as function of time and we compare different flavors of the overlap-based energy sharing, namely by combining with the fewest-switches surface hopping (FS) to select the active state (in light-brown and in yellow in the figure) and either by using Wigner-sampled initial conditions (in light-brown and orange), i.e., positions and momenta, or by initializing the momenta to same zero value while sampling only the positions (in yellow). In Fig. 5, we show as reference the results reported in Fig. 2 where a deterministic selection of the active state is done by imposing the trajectories to follow the electronic state with largest population (in orange). While, as expected, the population slightly changes by changing the initial conditions, the internal consistency is not altered since no frustrated hops all along the simulated dynamics are observed (in all cases the velocity is rescaled along the NACV after the hop and there are no frustrated hops).
In Fig. 5 (bottom panel), we also report the population of the diabatic state that is fully populated at the initial time. We remind that CCT-TSH calculations are performed in the adiabatic basis, however, the LVC model is constructed in the diabatic basis. The QuantumModelLib library provides the transformation matrix from diabatic-to-adiabatic basis and vice versa, which depends on the nuclear positions. Therefore, at all nuclear positions, we are capable of transforming the electronic coefficients yielding the representation of the electronic conditional amplitude in the adiabatic basis to the diabatic representation. Nonetheless, this transformation provides a meaningful estimate of the diabatic population only if the algorithm is internally consistent. This is indeed the case for the CCT-TSH results shown in Fig. 5. We compare the CCT-TSH diabatic populations to reference MCTDH results of Ref. [29]. The overlap-based scheme with deterministic selection of the active state shows an excellent agreement with the reference.
Another challenging test for the CCT-TSH algorithms is the S2-to-S1 photodynamics of DMABN. In this case, within the first 100 fs of dynamics, the system crosses multiple times nonadiabatic regions, such that the S2 state loses population almost completely in a step-wise manner to S1. We have recently shown 107 that DMABN stretches the approximations underlying CT-MQC and CT-TSH, but CCT-TSH successfully cures the issues of the original coupled-trajectory algorithms guaranteeing internal consistency. Below, we explore the robustness of various implementations of CCT-TSH in this respect.
In Fig. 6, we report the population of the state S2 as function of time, comparing CCT-TSH with TSH-ED. In all simulations, the velocity is rescaled along the NACV after the hop occurs and the velocity is inverted if a trajectory suffers from a frustrated hop. The population calculated using TSH-ED is shown in green in the figure, using dsahed lines for the fraction of trajectories (F) and continuous lines for the electronic populations (P). Note that in TSH-ED we observed a total of 529 frustrated hops. The equity-based (in yellow) and overlap-based (in orange) procedures for the energy sharing completely remove the occurrence of frustrated hops, yielding very similar populations, slightly larger than TSH-ED already after 500 a.t.u. The energy sharing with quantum-momentum-based scheme (in red) again produces 120 frustrated hops which does not seem to strongly affect the internal consistency. Nonetheless, the occurrence of frustrated hops suggests issues with the calculation of the quantum momentum, as observed previously for CT-MQC. 107
Figure 6: Adiabatic population decays of S2 in DMABN predicted by the dfferent surface hopping schemes, with TSH-ED (in green), with the equity-based scheme (in yellow), with the overlap-based scheme (in orange), with the quantum-momentum-based scheme (in red). Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
Finally, similarly to Fig. 5, we compare additional flavors of the overlap-based energy-sharing scheme, namely in combination with fewest switches and with partial sampling of the initial conditions. We report the population of S2 as function of time (top panel) and the population of the diabatic state that is almost completely populated at the initial time. We stress here that our trajectory-based calculations are initialized in the adiabatic state S2, which has a mixed diabatic character in the Franck-Condon region. This feature explains the rapid population decrease of S2 at very short times as well as the fact that the reported diabatic populations are not exactly one at the initial time. This difference in the initial condition might partly explain the deviation of the diabatic populations determined using CCT-TSH from the reference MCTDH results of Ref. [29]. Nevertheless, overall the agreement remains more than satisfactory.
Figure 7: Same as in Fig. 5 but for DMABN.
4 Conclusions